Schurz Corporation's management reports that its average delivery cycle time is 29 days, its average throughput time is 11.0 days, its manufacturing cycle efficiency (MCE) is 0.50, its average move time is 0.2 day, and its average queue time is 4.0 days.

Required: Round your answer to 1 decimal place.

What is the Wait Time?

Process Time?

Inspection Time?
Answer #1

Wait time = Delivery cycle time - Throughput time

= 29-11

= 18 days

Process time = Manufacturing cycle efficiency * Throughput time

= 0.50*11

= 5.5 days

Inspection time = Throughput time - process time - move time - queue time

= 11 - 5.5 - 0.2 - 4

= 1.3 days

Comment if you face any issues